OpenStudy (anonymous):

@ganeshie8 The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation? A. (x − 1)^2 − (y − 1)^2 = 1 B. (x − 1^)2 + (y + 1)^2 = 3 C. (x + 1)^2 + (y − 1)^2 = 1 D. (x − 1)^2 + (y + 1)^2 = 1

OpenStudy (anonymous):

Complete the squares

OpenStudy (anonymous):

\[ x^2 + y^2 − 2x + 2y − 1 = 0\\ x^2 − 2x+1-1+y^2 + 2y +1-1− 1 = 0\\ (x-1)^2+(y+1)^2-1-1-1=0\\ (x-1)^2+(y+1)^2 =3\\ \]

OpenStudy (anonymous):

B. is the answer

OpenStudy (anonymous):

your awesome in explaining it. Thanks :D

OpenStudy (anonymous):

YW